# Gravastar on the brane with a timelike extra dimension

### v1

@@ 3 Physical properties of the gravastar
3.1 Matter density and pressure of the shell
@@
$\rho=C_{1}e^{-\nu}={C_{1}}{{\rm e}^{-\frac{1}{2}{\frac{\left(\mu+1\right)B{r}^{2}}{\mu}}}}.$ (3.1)
@@
$\rho_{\mathrm{eff}}=8\pi\rho\left(r\right)\left[1-{\frac{\rho\left(r\right)}{\rho_{\mathrm{c}}}}\right]-{\frac{12U}{\rho_{\mathrm{c}}}}.$ (3.2)
@@
$p_{r}^{\mathrm{eff}}=\pi\left[p\left(r\right)-{\frac{\rho\left(r\right)\left(2p\left(r\right)+\rho\left(r\right)\right)}{\rho_{{c}}}}\right]-{\frac{4U}{\rho_{\mathrm{c}}}}-{\frac{8P}{\rho_{\mathrm{c}}}}.$ (3.3)
@@
$p_{t}^{\mathrm{eff}}=8\pi\left[\mu\rho\left(r\right)-{\frac{\rho\left(r\right)\left(2\mu\rho\left(r\right)+\rho\left(r\right)\right)}{\rho_{\mathrm{c}}}}\right]-{\frac{4U}{\rho_{\mathrm{c}}}}+{\frac{4P}{\rho_{\mathrm{c}}}}.$ (3.4)
Figure 1: Variation of the effective density w.r.t. the radial coordinate $r$.
@@ 3.2 Energy
$E$ $=$ $4\pi\int_{R}^{R+\epsilon}\rho^{\mathrm{eff}}r^{2}dr$ (3.5)
$=$ $4\pi{C_{1}}\left[-{\frac{r\mu}{\left(\mu+1\right)B}{{\rm e}^{-{\frac{\left(\mu+1\right)B{r}^{2}}{2\mu}}}}}+{\frac{\mu\sqrt{2\pi}}{2\left(\mu+1\right)B}{\rm erf}\left(\sqrt{{\frac{\left(\mu+1\right)Br}{2\mu}}}\right){\frac{1}{\sqrt{{\frac{\left(\mu+1\right)B}{\mu}}}}}}\right],$
where erferf denotes the error function.
Figure 2: Variation of the energy of the shell with respect to $r$.
@@
$E=4\pi{C_{1}}\left[-{\frac{r{{\rm e}^{-B{r}^{2}}}}{2B}}+{\frac{\sqrt{\pi}{\rm erf}\left(\sqrt{B}r\right)}{4{B}^{3/2}}}\right]$ (3.6)
@@ 3.3 Entropy
$S=\int_{R}^{R+\epsilon}4\pi r^{2}s(r)\sqrt{e^{\lambda}}dr,$ (3.7)
@@
$s(r)=\frac{\xi^{2}k_{B}^{2}T(r)}{4\pi\hbar^{2}}=\frac{\xi k_{B}}{\hbar}\sqrt{\frac{p}{2\pi}},$ (3.8)
@@
$S$ $=\frac{4\pi r^{2}\alpha k_{B}\epsilon}{\hbar}\Bigg[\frac{2Br^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}{\begin{aligned} &-24r^{2}C_{1}^{2}\pi e^{-2Br^{2}}+8\pi C_{1}r^{2}\rho_{\mathrm{c}}e^{-Br^{2}}\\ &\quad-8r^{2}\left(\omega+\frac{1}{2}\right)\sqrt{\rho_{0}C_{1}}\,e^{-\frac{1}{2}Br^{2}}+\rho_{\mathrm{c}}\end{aligned}}\Bigg]^{1/2}$ (3.10)
@@ 3.4 Proper thickness
$=\epsilon\Bigg[\frac{2Br^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}{\begin{aligned} &-24r^{2}C_{1}^{2}\pi\left(e^{-Br^{2}}\right)^{2}+8\pi C_{1}r^{2}\rho_{\mathrm{c}}e^{-Br^{2}}\\ &\quad-8r^{2}\left(\omega+\frac{1}{2}\right)\sqrt{\rho_{0}C_{1}}\,e^{-\frac{1}{2}Br^{2}}+\rho_{\mathrm{c}}\end{aligned}}\Bigg]^{1/2}$ (3.11)
The proper thickness quantifies the actual physical width of the gravastar shell as measured by a local observer, rather than a coordinate artifact. A finite, well-behaved thickness ensures that the shell represents a realistic matter layer capable of supporting stresses, storing energy and entropy, and smoothly mediating between the interior and exterior spacetimes. This confirms that the gravastar avoids unphysical surface singularities and provides a stable, physically meaningful alternative to an infinitesimally thin shell.
4 Boundary and matching conditions
@@
The appropriate framework for describing such junctions is provided by the Darmois–Israel formalism [60, 61], which relates the intrinsic stress–energy content of the shell to the discontinuity in the extrinsic curvature across the hypersurface. Within this approach, the surface stress–energy arises from the geometric mismatch between the interior and exterior spacetimes and encapsulates the physical properties of the shell.
As one approaches the outer boundary of the gravastar from within the shell, the variation of the energy distribution gives rise to a discontinuity in the extrinsic curvature. This geometric discontinuity manifests itself as intrinsic surface energy density and surface pressures on the junction hypersurface. The presence of this shell ensures a smooth matching of the interior and exterior regions in terms of the metric potentials, even though their derivatives may be discontinuous. Consequently, the gravastar remains geodesically complete and is supported by a well-defined matter configuration at the junction.
Following the prescription originally developed by Lanczos [62], the intrinsic surface stress–energy tensor $S^{i}_{\;j}$ associated with the shell is expressed as
$S^{i}_{\;j}=-\frac{1}{8\pi}\left(\kappa^{i}_{\;j}-\delta^{i}_{\;j}\kappa^{k}_{\;k}\right),$ (4.1)
@@
$\kappa_{ij}=\kappa^{+}_{ij}-\kappa^{-}_{ij}.$ (4.2)
@@
$\kappa^{\pm}_{ij}=-n^{\pm}_{\nu}\left[\frac{\partial^{2}X^{\nu}}{\partial\xi^{i}\partial\xi^{j}}+\Gamma^{\nu}_{\alpha\beta}\frac{\partial X^{\alpha}}{\partial\xi^{i}}\frac{\partial X^{\beta}}{\partial\xi^{j}}\right]_{S},$ (4.3)
@@
$n^{\pm}_{\nu}=\pm\left|g^{\alpha\beta}\frac{\partial f}{\partial X^{\alpha}}\frac{\partial f}{\partial X^{\beta}}\right|^{-\frac{1}{2}}\frac{\partial f}{\partial X^{\nu}},$ (4.4)
@@
$=-\frac{1}{4\pi R}\Bigg[\sqrt{\,1-\frac{2M}{R}-\frac{Q}{R^{2}}\,}-\sqrt{\,\frac{-8r^{2}\!\left(\omega+\tfrac{1}{2}\right)\sqrt{\rho_{0}C_{1}}+8\pi C_{1}\!\left(C_{1}-\rho_{c}\right)r^{2}+\rho_{c}}{2Br^{2}\rho_{c}+\rho_{c}}}\Bigg]$ (4.5)
@@
$\mathcal{P}$ $=$ $\frac{1}{16\pi}\bigg[\bigg(\frac{2-\lambda^{\prime}R}{R}\bigg)\sqrt{e^{-\lambda}}\bigg]_{-}^{+}$ (4.6)
$=$ $-{\frac{M-r}{{4r}^{2}\pi}{\frac{1}{\sqrt{1-{\frac{2M}{r}}-{\frac{Q}{{r}^{2}}}}}}}-{\frac{4\sqrt{3}\left(\frac{3}{2}{r}^{2}\sqrt{\rho_{{0}}{C_{1}}}+\pi{C_{1}}\left({C_{1}}-\rho_{{c}}\right){r}^{2}+\frac{3\rho_{\mathrm{c}}}{16}\right)}{3r\rho_{\mathrm{c}}\pi{\sqrt{{\frac{12{r}^{2}\sqrt{\rho_{{0}}{C_{1}}}+8\pi{C_{1}}\left({C_{1}}-\rho_{\mathrm{c}}\right){r}^{2}+3\rho_{\mathrm{c}}}{\rho_{\mathrm{c}}}}}}}}$
@@
$m_{s}$ $=$ $4\pi R^{2}\Sigma$ (4.7)
$=$ ${\frac{1}{4\pi r}\left(\sqrt{{\frac{8\pi{{C_{1}}}^{2}{r}^{3}-8\pi{C_{1}}{r}^{3}\rho_{\mathrm{c}}+12\sqrt{\rho_{{0}}{C_{1}}}{r}^{3}+3\rho_{\mathrm{c}}r}{3\rho_{\mathrm{c}}r}}}-\sqrt{1-{\frac{2M}{r}}-{\frac{Q}{{r}^{2}}}}\right)}$
@@
$M_{\rm Tot}$ $=\frac{1}{6\rho_{c}r}\Biggl[-12r^{4}\sqrt{\rho_{0}C_{1}}+8\pi r^{3}\Sigma\sqrt{3}\sqrt{\frac{12r^{2}\sqrt{\rho_{0}C_{1}}+8\pi C_{1}(C_{1}-\rho_{c})r^{2}+3\rho_{c}}{\rho_{c}}}\rho_{c}$
$\quad+\bigl((-48\pi^{2}\Sigma^{2}+8\pi C_{1})r^{4}-3Q\bigr)\rho_{c}-8\pi C_{1}^{2}r^{4}\Biggr].$ (4.8)
@@
$1-\frac{2M}{R_{2}}-\frac{Q}{R_{2}^{2}}=C^{2}e^{BR_{2}^{2}},$ (4.9)
$\frac{2M}{R_{2}^{2}}+\frac{2Q}{R_{2}^{3}}=2BR_{2}C^{2}e^{BR_{2}^{2}}.$ (4.10)
Similarly, the radial metric component $g_{rr}$ has been matched with the interior metric at $r=R_{1}$ and with the exterior metric at $r=R_{2}$, ensuring a smooth interpolation of the geometry across the shell. These matching conditions guarantee that the shell acts as a proper junction between the interior and exterior spacetimes while maintaining geodesic completeness and regularity of the metric.
Using the chosen model parameters, $M=3.35M_{\odot}$, $R_{1}=10km$, and $R_{2}=10.3km$, we solve the above matching conditions to determine the remaining constants of the system. The resulting values are:
$Q$ $=0.08865696775,$ $B$ $=0.008803696661km^{-2},$
$\omega$ $=0.0013907304225,$ $C_{1}$ $=0.04236407711,$
$\rho_{0}$ $=9.807751458\times 10^{-7}m^{4},$ $\rho_{c}$ $=0.05117029034m^{4}.$
These constants fully specify the gravastar model, linking the interior de Sitter-like core, the finite shell, and the exterior vacuum with tidal charge. The smooth matching ensures both the continuity of the metric potentials and a physically consistent distribution of matter and stress across the shell, while the derivative matching guarantees that the gravitational field behaves regularly at the outer boundary. The numerically determined values of the model parameters are physically well justified. The relatively small tidal charge indicates that even slight deviations from a Schwarzschild vacuum, induced by bulk gravitational effects projected onto the brane through the Weyl tensor, can lead to noticeable modifications of the braneworld dynamics. Likewise, the small magnitude of the equation-of-state parameter $\omega$ suggests that the effective fluid representing the bulk contributions on the brane behaves consistently with standard energy conditions and can effectively mimic additional ordinary matter on the brane. Furthermore, the critical energy density on the brane, influenced by the brane tension $\sigma$, is found to be small. This highlights that braneworld corrections play a crucial role in shaping the gravastar configuration, placing the system in a regime where standard general relativity is significantly modified. This contrasts with the large $\sigma$ limit, in which general relativity is effectively recovered.
5 Stability

### v2

@@ III Physical properties of the gravastar
III.1 Matter density and pressure of the shell
@@
$\rho=C_{1}e^{-\nu}=C_{1}\exp\left[-\frac{1}{2}\frac{(\mu+1)Br^{2}}{\mu}\right].$ (31)
@@
$\rho_{\mathrm{eff}}=8\pi\rho\left(r\right)\left[1-{\frac{\rho\left(r\right)}{\rho_{\mathrm{c}}}}\right]-{\frac{12U}{\rho_{\mathrm{c}}}}.$ (32)
@@
$p_{r}^{\mathrm{eff}}=\pi\left[p\left(r\right)-{\frac{\rho\left(r\right)\left(2p\left(r\right)+\rho\left(r\right)\right)}{\rho_{{c}}}}\right]-{\frac{4U}{\rho_{\mathrm{c}}}}-{\frac{8P}{\rho_{\mathrm{c}}}}.$ (33)
@@
$p_{t}^{\mathrm{eff}}=8\pi\left[\mu\rho\left(r\right)-{\frac{\rho\left(r\right)\left(2\mu\rho\left(r\right)+\rho\left(r\right)\right)}{\rho_{\mathrm{c}}}}\right]-{\frac{4U}{\rho_{\mathrm{c}}}}+{\frac{4P}{\rho_{\mathrm{c}}}}.$ (34)
Figure 1: Variation of the effective energy density of the gravastar shell, $\rho_{\mathrm{eff}}$, as a function of the radial coordinate $r$ within the finite shell region $R\leq r\leq R+\epsilon$. The different curves represents the profile of $\rho_{\mathrm{eff}}(r)$ obtained from Eq. (32) for different masses of the gravastar. The numerical evaluation is performed using the model parameters $B=0.011415~{\rm km^{-2}}$, $Q=0.13822104$, $\rho_{0}=0.05\times 10^{-7}~{\rm m}^{4}$ together with the constants determined from the matching conditions given in Table I. The plot shows that the effective density attains its maximum value at the inner boundary of the shell and decreases monotonically as a function of $r$ toward the outer boundary, indicating a smooth and regular matter distribution across the shell.
@@ III.2 Energy
$E$ $=$ $4\pi\int_{R}^{R+\epsilon}\rho^{\mathrm{eff}}r^{2}dr$ (35)
$=$ $4\pi{C_{1}}\left[-{\frac{r\mu}{\left(\mu+1\right)B}{{e}^{-{\frac{\left(\mu+1\right)B{r}^{2}}{2\mu}}}}}+{\frac{\mu\sqrt{2\pi}}{2\left(\mu+1\right)B}{\text{erf}}\left(\sqrt{{\frac{\left(\mu+1\right)Br}{2\mu}}}\right){\frac{1}{\sqrt{{\frac{\left(\mu+1\right)B}{\mu}}}}}}\right],$
where erf denotes the error function.
Figure 2: Energy contained within the gravastar shell, $E$, as a function of the radial coordinate $r$ obtained from Eq. (36). The different curves corresponding to different masses of the gravastar show the radial dependence of the shell energy for the parameter values $B=0.011415~{\rm km^{-2}}$, $Q=0.13822104$ and $\rho_{0}=0.05\times 10^{-7}~{\rm m}^{4}$ with the remaining constants fixed by the boundary conditions. The energy increases monotonically with increasing $r$ from the interior boundary $r=R$ to the exterior boundary $r=R+\epsilon$. This behaviour reflects the gradual accumulation of effective matter energy throughout the finite shell region.
@@
$E=4\pi{C_{1}}\left[-{\frac{r{{e}^{-B{r}^{2}}}}{2B}}+{\frac{\sqrt{\pi}{\text{erf}}\left(\sqrt{B}r\right)}{4{B}^{3/2}}}\right].$ (36)
@@ III.3 Entropy
$S=\int_{R}^{R+\epsilon}4\pi r^{2}s(r)\sqrt{e^{\lambda}}dr,$ (37)
@@
$s(r)=\frac{\xi^{2}k_{B}^{2}T(r)}{4\pi\hbar^{2}}=\frac{\xi k_{B}}{\hbar}\sqrt{\frac{p}{2\pi}},$ (38)
@@
$S$ $=\frac{4\pi r^{2}\alpha k_{B}\epsilon}{\hbar}\left[\frac{2Br^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}{-24r^{2}C_{1}^{2}\pi e^{-2Br^{2}}+8\pi C_{1}r^{2}\rho_{\mathrm{c}}e^{-Br^{2}}-8r^{2}\left(w+\frac{1}{2}\right)\sqrt{\rho_{0}C_{1}}e^{-\frac{1}{2}Br^{2}}+\rho_{\mathrm{c}}}\right]^{1/2}.$ (40)
@@ III.4 Proper thickness
$=\epsilon\Bigg[\frac{2Br^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}{\begin{aligned} &-24r^{2}C_{1}^{2}\pi e^{-2Br^{2}}+8\pi C_{1}r^{2}\rho_{\mathrm{c}}e^{-Br^{2}}-4r^{2}\left(2w+1\right)\sqrt{\rho_{0}C_{1}}e^{-\frac{1}{2}Br^{2}}+\rho_{\mathrm{c}}\end{aligned}}\Bigg]^{1/2}.$ (41)
Figure 3: Proper thickness of the gravastar shell, $\ell$, as a function of the radial coordinate $r$ within the thin shell region $R\leq r\leq R+\epsilon$. The curves represent the numerical evaluation of the proper thickness for different masses of the gravastar obtained from Eq. (34). The calculation is performed using the parameter values $B=0.011415~{\rm km^{-2}}$, $Q=0.13822104$ and $\rho_{0}=0.05\times 10^{-7}~{\rm m}^{4}$, together with the constants determined from the matching conditions at the shell boundaries. The plot shows that the proper thickness decreases smoothly as a function of $r$ across the shell, indicating that the shell region remains finite and well behaved throughout the gravastar configuration.
The proper thickness quantifies the actual physical width of the gravastar shell as measured by a local observer, rather than a coordinate artifact. A finite, well-behaved thickness ensures that the shell represents a realistic matter layer capable of supporting stresses, storing energy and entropy, and smoothly mediating between the interior and exterior spacetimes. This confirms that the gravastar avoids unphysical surface singularities and provides a stable, physically meaningful alternative to an infinitesimally thin shell. Fig. 3 depicts the behavior of the proper thickness of the gravastar shell as a function of the radial coordinate within the shell region. The plot shows that the proper thickness decreases gradually with increasing radial distance, indicating that the shell becomes progressively thinner toward the outer boundary. This behavior is consistent with the thin-shell structure of the gravastar and ensures a smooth transition between the interior de Sitter-like core and the exterior braneworld geometry.
IV Boundary and matching conditions
@@
The appropriate framework for describing such junctions is provided by the Darmois–Israel formalism Darmois ; Israel , which relates the intrinsic stress–energy content of the shell to the discontinuity in the extrinsic curvature across the hypersurface. Within this approach, the surface stress–energy arises from the geometric mismatch between the interior and exterior spacetimes and encapsulates the physical properties of the shell.
As one approaches the outer boundary of the gravastar from within the shell, the variation of the energy distribution gives rise to a discontinuity in the extrinsic curvature. This geometric discontinuity manifests itself as an intrinsic surface energy density and surface pressures on the junction hypersurface. The presence of this shell ensures smooth matching of the interior and exterior regions in terms of the metric potentials, although their derivatives may be discontinuous. Consequently, the gravastar remains geodesically complete and is supported by a well-defined matter configuration at the junction at $r=R$.
Following the prescription originally developed by Lanczos Lanczos , the intrinsic surface stress–energy tensor $S^{i}_{\;j}$ associated with the shell is expressed as
$S^{i}_{\;j}=-\frac{1}{8\pi}\left(\kappa^{i}_{\;j}-\delta^{i}_{\;j}\kappa^{k}_{\;k}\right),$ (42)
@@
$\kappa_{ij}=\kappa^{+}_{ij}-\kappa^{-}_{ij}.$ (43)
@@
$\kappa^{\pm}_{ij}=-n^{\pm}_{\nu}\left[\frac{\partial^{2}X^{\nu}}{\partial\xi^{i}\partial\xi^{j}}+\Gamma^{\nu}_{\alpha\beta}\frac{\partial X^{\alpha}}{\partial\xi^{i}}\frac{\partial X^{\beta}}{\partial\xi^{j}}\right]_{S},$ (44)
Table 1: Variation of different model parameters with gravastar mass.
Mass $R_{1}$ $R_{2}$ Thickness $(\epsilon)$ $w$ $\rho_{\mathrm{c}}$ $C$ $C_{1}$
2.50 8.15 8.2535 0.1035 0.0000769 1.5088018 0.4244991 1.0574555
2.75 8.57 8.7100 0.1400 0.2674440 1.5452259 0.3927440 1.1626156
3.00 9.00 9.1625 0.1625 0.4930296 1.7418696 0.3629314 1.4309239
3.25 9.40 9.6106 0.2106 0.9256219 2.8489255 0.3350389 2.6239795
@@
$n^{\pm}_{\nu}=\pm\left|g^{\alpha\beta}\frac{\partial f}{\partial X^{\alpha}}\frac{\partial f}{\partial X^{\beta}}\right|^{-\frac{1}{2}}\frac{\partial f}{\partial X^{\nu}},$ (45)
@@
$=-\frac{1}{4\pi R}\Bigg[\sqrt{1-\frac{2M}{R}-\frac{Q}{R^{2}}}-\sqrt{\frac{-4r^{2}\!\left(2w+1\right)\sqrt{\rho_{0}C_{1}}+8\pi C_{1}\!\left(C_{1}-\rho_{\mathrm{c}}\right)r^{2}+\rho_{\mathrm{c}}}{2Br^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}}\Bigg].$ (46)
@@
$\mathcal{P}$ $=$ $\frac{1}{16\pi}\bigg[\bigg(\frac{2-\lambda^{\prime}R}{R}\bigg)\sqrt{e^{-\lambda}}\bigg]_{-}^{+}$ (47)
$=$ ${\frac{R-M}{4\pi R^{2}}{\frac{1}{\sqrt{1-{\frac{2M}{R}}-{\frac{Q}{{R}^{2}}}}}}}-{\frac{4\sqrt{3}\left(24{R}^{2}\sqrt{\rho_{{0}}{C_{1}}}+16\pi{C_{1}}\left({C_{1}}-\rho_{{c}}\right){R}^{2}+3\rho_{\mathrm{c}}\right)}{3\pi R{\sqrt{\rho_{\mathrm{c}}\bigg[{12{r}^{2}\sqrt{\rho_{{0}}{C_{1}}}+8\pi{C_{1}}\left({C_{1}}-\rho_{\mathrm{c}}\right){r}^{2}+3\rho_{\mathrm{c}}\bigg]}}}}}.$
@@
$\hskip-14.22636ptm_{\mathrm{s}}$ $=$ $4\pi R^{2}\Sigma$ (48)
$=$ ${\frac{1}{4\pi R}\left(\sqrt{{\frac{8\pi{{C_{1}}}^{2}{R}^{3}-8\pi{C_{1}}{R}^{3}\rho_{\mathrm{c}}+12\sqrt{\rho_{{0}}{C_{1}}}{R}^{3}+3\rho_{\mathrm{c}}R}{3\rho_{\mathrm{c}}R}}}-\sqrt{1-{\frac{2M}{R}}-{\frac{Q}{{R}^{2}}}}\right)}.$
Table I summarizes the numerical values of the principal physical parameters characterizing the gravastar configuration for different choices of the total mass of the object. For each mass value, the corresponding inner and outer radii of the shell, $R_{1}$ and $R_{2}$, are obtained from the junction conditions, from which the shell thickness $\epsilon=R_{2}-R_{1}$ is determined. It is observed that as the gravastar mass increases, both the inner and outer radii increase accordingly, indicating that more massive configurations correspond to larger compact objects. Consequently, the shell thickness also increases gradually, which is physically reasonable since a larger gravitational mass requires a wider transition region connecting the interior de Sitter–like core to the exterior braneworld vacuum geometry.
The table further shows that the critical energy density $\rho_{\mathrm{c}}$ increases with the mass of the configuration. In the braneworld framework, $\rho_{\mathrm{c}}$ is directly related to the brane tension and controls the strength of the high-energy corrections to the effective gravitational field equations. Larger values of $\rho_{\mathrm{c}}$ correspond to weaker brane corrections, and in the limit $\rho_{\mathrm{c}}\rightarrow\infty$ the quadratic matter terms vanish, thereby recovering standard General Relativity. Furthermore, the equation-of-state parameter $w$ also increases gradually with the mass. Since $w$ characterizes the thermodynamic behavior of the shell matter, its increase reflects a modification of the pressure–energy density relation required to sustain equilibrium in more massive gravastar configurations. This behaviour ensures that the shell matter remains sufficiently stiff to support the structure against gravitational collapse while maintaining the smooth matching between the interior and exterior geometries. The integration constants $C$ and $C_{1}$ listed in the table are determined from the matching conditions and adjust consistently with the other parameters so that the metric functions and physical quantities remain regular throughout the configuration. These trends collectively indicate that the obtained parameter values lead to a physically consistent and stable gravastar structure within the braneworld framework.
@@
$M_{\rm Tot}$ $=\frac{1}{6\rho_{\mathrm{c}}R}\Biggl[8\pi R^{3}\Sigma\sqrt{3}\sqrt{\rho_{\mathrm{c}}\left(12R^{2}\sqrt{\rho_{0}C_{1}}+8\pi C_{1}(C_{1}-\rho_{\mathrm{c}})R^{2}+3\rho_{\mathrm{c}}\right)}-12R^{4}\sqrt{\rho_{0}C_{1}}$
$\quad+\bigg((8\pi C_{1}-48\pi^{2}\Sigma^{2})R^{4}-3Q\bigg)\rho_{\mathrm{c}}-8\pi C_{1}^{2}R^{4}\Biggr].$ (49)
@@
$1-\frac{2M}{R_{2}}-\frac{Q}{R_{2}^{2}}=C^{2}e^{BR_{2}^{2}},$ (50)
$\frac{2M}{R_{2}^{2}}+\frac{2Q}{R_{2}^{3}}=2BR_{2}C^{2}e^{BR_{2}^{2}}.$ (51)
Similarly, the radial metric component $g_{rr}$ has been matched with the interior metric at $r=R_{1}$ and with the exterior metric at $r=R_{2}$, ensuring a smooth interpolation of the geometry across the shell. These matching conditions guarantee that the shell acts as a proper junction between the interior and exterior spacetimes while maintaining geodesic completeness and regularity of the metric. Using the chosen model parameters, $Q=0.138221$, $B=0.011415$ km-2, $\rho_{0}=0.05\times 10^{-7}$ m4. We solve the above matching conditions to determine the remaining constants of the system.
These constants fully specify the gravastar model, linking the interior de Sitter-like core, the finite shell, and the exterior vacuum with tidal charge. The smooth matching ensures both the continuity of the metric potentials and a physically consistent distribution of matter and stress across the shell, while the derivative matching guarantees that the gravitational field behaves regularly at the outer boundary. The numerically determined values of the model parameters are physically well justified. The relatively small tidal charge indicates that even slight deviations from the Schwarzschild vacuum, induced by bulk gravitational effects projected onto the brane through the Weyl tensor, can lead to noticeable modifications of the braneworld dynamics. Likewise, the small magnitude of the equation-of-state parameter $w$ suggests that the effective fluid representing the bulk contributions on the brane behaves consistently with standard energy conditions and can effectively mimic additional ordinary matter on the brane. Furthermore, the critical energy density on the brane, influenced by the brane tension $\sigma$, is found to be small. This highlights that braneworld corrections play a crucial role in shaping the gravastar configuration, placing the system in a regime where standard general relativity is significantly modified. This contrasts with the large $\sigma$ limit, in which general relativity is effectively recovered.
V Stability